OpenStudy (anonymous):

Can you check these for me? Thanks! Determine the type of boundary line and shading for the graph of the inequality -5x - y less than or equal to -2 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin. I chose C. Which of the following ordered pairs is a solution to the inequality y less than or equal to -2x + 10? (0, -4) (1, 13) (6, -1) (3, 8)

OpenStudy (anonymous):

I chose B

OpenStudy (jim766):

1st one -5x - y <= -2 add 5x to both sides + 5x +5x - y <= 5x -2 divide by neg 1, flip the sign y >= -5x + 2 is >= a dotted or solid line?

OpenStudy (anonymous):

Uhm dotted?

OpenStudy (jim766):

greater than equal is solid, whenever it is equal, it is solid shaded above or below?

OpenStudy (anonymous):

Above?

OpenStudy (jim766):

right, it will look like this:  shaded above, which is correct?

OpenStudy (anonymous):

I don't know. I understand it kinda but I really just don't get it at all.

OpenStudy (jim766):

what is the origin on a graph?

OpenStudy (jim766):



OpenStudy (anonymous):

(0,0)??

OpenStudy (jim766):

right , see the picture above is the origin shaded?

OpenStudy (anonymous):

yes i think

OpenStudy (jim766):

yes it is,,,,so solid line with origin shaded

OpenStudy (jim766):

y <= -2x + 10 to do this one substitute in the x and y from the points the gave you ( x , Y) ie -4 <= -2(0) + 10 -4 <= 0 + 10 -4 <= 10 this is true, so it is in the solution

OpenStudy (jim766):

13 <= -2(1) + 10 13< = -2 + 10 13 < = 8 not true, not in the solution

OpenStudy (jim766):

make sense?

OpenStudy (anonymous):

Not really im confused

OpenStudy (jim766):

do you see the points they want you to choose from? They want us to see if they are in the shaded area of the graph or not. The easiest way to do this to substitute each in the inequality and see if it is true or not

OpenStudy (jim766):

look the first point the give us to try x y (0, -4) just substitute these in -4 <= -2(0) + 10 -4 <= 0 + 10 -4 <= 10 this is true, so it is in the solution -4 is less than or equal to 10, so it is true , make sense?

OpenStudy (anonymous):

I guess it kinda does.

OpenStudy (jim766):

look at the 2nd one x y y <= -2x + 10 (1, 13) 13 <= -2(1) + 10 13< = -2 + 10 13 < = 8 not true, not in the solution 13 is not less than or equal to 8, it is greater, so not true.. my guess is that you wont have to check the others A is correct

OpenStudy (anonymous):

I think I get it but I will probably forget it by tomorrow

OpenStudy (jim766):

I know the feeling....hang in there...